OpenStudy (anonymous):

What is the slope intercept form of the equation 3x-y= - 6?

OpenStudy (anonymous):

well what is slope intercept form?

OpenStudy (anonymous):

y=mx + c

OpenStudy (anonymous):

^ lol

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

y = zx + e :P

OpenStudy (anonymous):

3x-y= - 6 3x + 6 = y

OpenStudy (anonymous):

@Yahoo! they're supposed to do it lol

OpenStudy (anonymous):

y=3x + 6 :D

OpenStudy (anonymous):

lol yes :P

OpenStudy (anonymous):

Thanks!